OpenStudy (anonymous):

solve for the variable a-(-17)=-30 so you add 17 both side or what?

OpenStudy (anonymous):

a-(-17)=-30 a+17=-30 a=-30-17 a=-47

OpenStudy (anonymous):

thanks